integrate the following integral
∫ 1 / 1 + cos 3x dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int \frac{1}{1+cos3x}dx
I = \int \frac{1}{2cos^{2}\frac{3x}{2}}dx
I = \int \frac{sec^{2}\frac{3x}{2}}{2}dx
I = \frac{tan\frac{3x}{2}}{2.\frac{3}{2}}+c
I = \frac{tan\frac{3x}{2}}{3}+c